There are six multiple-choice questions on an exam, each with three possible answers. Determine the number of possible answer sequences for the six questions
Found 2 solutions by stanbon, drk:
Answer by stanbon(57294)

There are six multiple-choice questions on an exam, each with three possible answers. Determine the number of possible answer sequences for the six questions
------------
1st question: 3 choices
2nd question: 3 choices
---
---
6th question: 3 choices
=============================
Total # of possible sequences = 3^6 = 729
==============================================
Cheers,
Stan H.

Answer by drk(1908)

It will be 3^6 = 729
---
it is always (choices per question)^ (number of questions)